# Real-Time Summarization of Twitter

### v1

@@ II-B Data and Preprocessing
In order to evaluate our filtering system in the real time environment, we need the queries (interest profiles), a stream of tweets and the corresponding ground truth [17].
A collection of the interest profiles (queries) are provided by TREC 2016  [^1] http://trecrts.github.io/TREC2016-RTS-topics.json and we used the narrative part in the interest profile for the queries.
@@ III-B2 Future Work
  • Scalability: Address the challenge of scaling up for higher tweet frequencies or larger user bases by exploring distributed computing frameworks such as Apache Spark and optimized data structures for fast retrieval. Implementing real-time processing pipelines and evaluating system performance under varying tweet volumes and different user interest profiles will also be crucial.

### v2

@@ II-B Data and Preprocessing
In order to evaluate our filtering system in the real time environment, we need the queries (interest profiles), a stream of tweets and the corresponding ground truth [17, 19, 32].
A collection of the interest profiles (queries) are provided by TREC 2016  [^1] http://trecrts.github.io/TREC2016-RTS-topics.json and we used the narrative part in the interest profile for the queries [21, 23, 35, 33].
@@ III-B2 Future Work
  • Scalability: Address the challenge of scaling up for higher tweet frequencies or larger user bases by exploring distributed computing frameworks such as Apache Spark and optimized data structures for fast retrieval. Implementing real-time processing pipelines and evaluating system performance under varying tweet volumes and different user interest profiles will also be crucial [36].